Suppose two distinct integers are chosen from between 5 and 17, inclusive. What is the probability that their product is odd?
Found 2 solutions by solver91311, sudhanshu_kmr:
Answer by solver91311(24713)



In order for a product to be odd, both factors must be odd. That is because by definition a number with an even factor is even.

There are 13 numbers from 5 to 17 inclusive. 7 of them are odd. There is a probability that the first number selected will be odd. Given that the first number selected is odd, there is a probability that the next number will be odd given that the second number selected has to be different from the first.



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by sudhanshu_kmr(1152)


there are 13 integers, 7 are odds and 6 are evens....
product of two integers will be odd if both are odd integers...

no. of ways to choose 2 odd integer out of 7 = 7C2 = 21

no. of ways to choose 2 integers out of 13 integers = 13C2 = 78


probability = 21/78 = 7/26